# DGFusion: Depth-Guided Sensor Fusion for Robust Semantic Perception

### v2

@@ III-C Implementation Details
We use a Swin-T backbone [liu2021swin] pre-trained on ImageNet [deng2009imagenet] and follow the training protocol of CAFuser [cafuser]. DGFusion is trained with a batch size of 8 for 180k iterations on MUSES and 200k on DeLiVER. For the depth branch, we set $\tau=0.8$, $\lambda_{\text{L1}}=0.9$, $\lambda_{\text{es}}=0.05$, $\lambda_{\text{pes}}=0.05$, and $\lambda_{\text{depth}}=1$. We use the OneFormer head [jain2023oneformer] for our semantic branch, which enables semantic, instance, and panoptic segmentation via a single model, and we adopt the default settings for the segmentation and CT losses from OneFormer and CAFuser, respectively. All input modalities are normalized over the entire dataset and the data augmentations from CAFuser are applied. Our network is optimized using AdamW with a poly learning rate scheduler.
@@
Mask2Former [cheng2022masked] C Swin-T 70.7
SegFormer [xie2021segformer] C MiT-B2 72.5
OneFormer [jain2023oneformer] C Swin-T 72.8
CMNeXt [zhang2023delivering] CLRE MiT-B2 72.1
GeminiFusion [jia2024geminifusion] CLRE MiT-B2 75.3
CAFuser-CAA [cafuser] CLRE Swin-T 78.5
CAFuser [cafuser] CLRE Swin-T 78.2
@@ IV Experiments
We evaluate DGFusion on two challenging multi-sensor datasets with a focus on adverse conditions in driving scenes: MUSES and DeLiVER. MUSES [MUSES] is a real-world dataset that comprises 2500 multi-sensor scenes captured under eight environmental conditions (rain, snow, fog, and clear weather, each in day/night) and containing high-quality human-annotated panoptic and semantic ground truth. Each sample includes RGB camera, lidar, radar, and event data, all of which we project onto the RGB plane with the help of the official SDK. We perform all our ablations on this challenging real-world dataset. Due to computational constraints, the hyperparameter ablation in Tab. VI was performed on half of the training data. DeLiVER [zhang2023delivering] is a synthetic dataset, featuring 7885 scenes recorded under adverse conditions (cloudy, foggy, night, rainy) and corner-case sensor failures (e.g. motion blur, lidar jitter) with semantic ground truth. Each scene provides RGB camera, lidar, and event camera modalities as well as a pixel-perfect depth map of the scenes. As this depth map is not a direct sensor recording and not realistic to obtain in this quality in any real-world scenario, we argue that one should treat it rather as an additional ground truth. Hence, we use the depth map as the dense supervision for the depth head of our method. As previous methods treat it as an input, we compare to such methods on both scenarios, with (CLDE) and without (CLE) the depth map as an input.
@@ IV-A Comparison with The State of The Art
DGFusion sets the new state of the art (SOTA) in multi-sensor panoptic and semantic segmentation compared to prior methods. As presented in Tab. I, we outperform previous methods on MUSES across all metrics, including a variety of adverse weather and lighting conditions. We set the new SOTA with 61.03% PQ (+1.33%) and improvements are most pronounced in adverse conditions. Compared with the previous SOTA method CAFuser [cafuser], all adverse conditions show a substantial improvement: 1.34% PQ in Fog, 1.63% PQ in Rain and even +2.57% PQ in Snow conditions, compared to a moderate +0.8% PQ in Clear conditions. This confirms our hypothesis that fusing both global condition cues and local depth features is crucial for robust performance under adverse environmental conditions.
@@
CMNeXt [zhang2023delivering] MiT-B2 50.3 53.0
StitchFusion [li2024stitchfusion] MiT-B2 50.8 53.4
GeminiFusion [jia2024geminifusion] MiT-B2 50.5 54.5
CAFuser-CAA [cafuser] Swin-T 51.2 55.2
CAFuser [cafuser] Swin-T 51.3 55.6
@@
As the OneFormer segmentation head can predict both panoptic and semantic segmentation, we also compare DGFusion to SOTA multimodal semantic segmentation methods on the MUSES dataset. Tab. II shows that our method significantly outperforms both camera-only baselines and prior multimodal approaches, achieving 79.5% mIoU (+ 1.0%). Notably, we surpass other fusion-based methods, such as CMNeXt [zhang2023delivering] and CAFuser [cafuser], highlighting the effectiveness of our multi-task setup for improving semantic segmentation in adverse-weather driving scenes.
@@
MUSES [MUSES]
OneFormer [jain2023oneformer]
CAFuser [cafuser]
@@ IV-C Qualitative Results
In Fig. 3, we present qualitative prediction results on MUSES under various conditions, demonstrating the strengths of our approach compared to existing methods. Under the clear day scenario, we accurately capture the tram in the distance and all four pedestrians on the left sidewalk, whereas other baselines overlook some of these persons. In the second example, our model successfully segments the person on the right edge and the cyclist on the left and even detects one of the smaller signs in the background. In the foggy night scene, we are the only method that detects both the cyclist and the bicycle despite the very low visibility. Finally, in the snow daytime scene, DGFusion is the only method that predicts at least one of the two pedestrians on the left sidewalk, though it misses the second person, highlighting a limitation in detecting heavily occluded objects. Our method also occasionally confuses similar classes, such as “wall” and “fence” (last row). Nevertheless, these examples highlight how our depth-guided method improves detection under all environmental conditions.
@@ V Conclusion
We have presented DGFusion, a novel depth-guided sensor fusion network for robust semantic perception. Our approach reformulates the multi-sensor fusion task as a multi-task problem by integrating an auxiliary depth head and directly leveraging noisy lidar input as additional supervision. In addition, we propose a depth-guided fusion module that combines local depth tokens with global condition tokens to adaptively guide sensor fusion in any condition. Extensive evaluations on both the real-world MUSES dataset and the synthetic DeLiVER dataset show that DGFusion outperforms previous state-of-the-art methods in panoptic and semantic segmentation, particularly in adverse conditions such as night, fog, rain, and snow. Moreover, our ablation studies confirm the essential contribution of our multi-task loss design and robust depth supervision, demonstrating the value of leveraging noisy lidar depth signals without incurring extra annotation costs. These advancements position DGFusion as a promising solution for enhancing semantic perception in autonomous driving and robotics under diverse and challenging environmental scenarios.
TABLE VIII: Class-wise PQ results on the MUSES test split.
Method
road
sidewalk
building
wall
fence
pole
traffic light
traffic sign
vegetation
terrain
sky
person
rider
car
truck
bus
train
motorcycle
bicycle
Mask2Former 95.12 72.08 76.59 38.35 28.96 39.85 37.54 48.73 69.73 49.60 83.95 34.03 19.31 55.70 28.59 42.58 38.21 14.88 17.06
MaskDINO 95.22 69.06 77.87 31.32 32.81 44.31 37.52 44.56 70.81 42.23 84.51 46.88 27.65 62.22 35.41 41.41 38.07 30.89 26.61
OneFormer 95.95 75.68 81.29 44.13 37.30 47.80 48.21 56.87 72.37 52.53 88.50 47.07 29.92 64.41 40.94 62.03 47.37 31.14 25.44
HRFuser 94.35 69.15 71.73 34.71 24.43 35.61 30.04 51.23 67.62 46.49 82.16 33.50 18.89 56.49 26.87 31.53 29.73 19.94 9.70
MUSES 95.86 76.42 80.40 47.43 37.56 45.57 42.83 60.23 75.01 53.83 83.69 42.21 33.51 63.65 40.01 45.09 42.65 28.52 23.88
CAFuser-CAA 96.14 78.29 83.12 51.40 42.48 49.35 53.79 64.07 78.65 56.09 90.65 51.79 39.53 66.33 46.88 58.99 52.51 39.07 29.18
CAFuser 96.36 79.52 84.38 50.87 43.55 49.84 50.82 63.18 79.40 56.90 92.11 50.51 39.05 65.69 45.95 62.23 54.34 40.63 28.96
DGFusion (Ours) 96.63 80.80 84.51 53.60 43.56 52.54 54.39 65.56 79.41 57.22 93.63 51.74 42.17 67.13 49.33 64.07 53.26 39.41 28.36
Detailed Class-Level Results
We present detailed class-level panoptic quality results on the MUSES test split in Table VIII. Our method consistently achieves superior performance across most classes. This highlights our model’s robust adaptation across varied semantic categories compared to other strong multimodal fusion methods, reflecting the effectiveness of our depth-guided fusion strategy.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
RGB
Lidar
ID 1 of Tab. 5 ($\lambda_{\text{es}}=0;\lambda_{\text{pes}}=0;\tau=1$)
DGFusion-depth (Ours)
Figure 4: Visual study Qualitative ablation on our loss design on the MUSES dataset. Best viewed on screen at full zoom.
Visual Ablation on Our Loss Design
In  Figure 4, we visually explore the effect of our loss design, which we investigate quantitatively in the main paper. The results illustrate that our proposed loss enhances the depth estimation quality by targeting the challenges of using noisy lidar as ground truth. Specifically, structural artifacts such as holes in building facades are effectively reduced, and boundaries are notably cleaner. Thereby demonstrating the strength of our method in handling challenging visual scenarios and providing visually superior depth predictions.
Further Qualitative Results
In Figure 5, we present further qualitative comparisons across diverse scenarios on the MUSES dataset. Our proposed method DGFusion shows improved robustness in scenarios challenging for RGB-only methods and baseline fusion approaches, such as perception at night, fog, and detection of distant objects. Specifically, it achieves superior instance separation and effectively leverages multimodal information under adverse conditions, thereby visually confirming the quantitative improvements in panoptic segmentation accuracy.
RGB
Lidar
Radar
Events
Ours-depth
GT
MUSES
OneFormer
CAFuser
DGFusion (Ours)
Figure 5: Further qualitative results on MUSES. Best viewed on a screen at full zoom.

### v3

@@ III-C Implementation Details
We use a Swin-T backbone [23] pre-trained on ImageNet [11] and follow the training protocol of CAFuser [4]. DGFusion is trained with a batch size of 8 for 180k iterations on MUSES and 200k on DeLiVER. For the depth branch, we set $\tau=0.8$, $\lambda_{\text{L1}}=0.9$, $\lambda_{\text{es}}=0.05$, $\lambda_{\text{pes}}=0.05$, and $\lambda_{\text{depth}}=1$. We use the OneFormer head [17] for our semantic branch, which enables semantic, instance, and panoptic segmentation via a single model, and we adopt the default settings for the segmentation and CT losses from OneFormer and CAFuser, respectively. All input modalities are normalized over the entire dataset and the data augmentations from CAFuser are applied. Our network is optimized using AdamW with a poly learning rate scheduler.
@@
Mask2Former [9] C Swin-T 70.7
SegFormer [28] C MiT-B2 72.5
OneFormer [17] C Swin-T 72.8
CMNeXt [29] CLRE MiT-B2 72.1
GeminiFusion [18] CLRE MiT-B2 75.3
CAFuser-CAA [4] CLRE Swin-T 78.5
CAFuser [4] CLRE Swin-T 78.2
@@ IV Experiments
We evaluate DGFusion on two challenging multi-sensor datasets with a focus on adverse conditions in driving scenes: MUSES and DeLiVER. MUSES [2] is a real-world dataset that comprises 2500 multi-sensor scenes captured under eight environmental conditions (rain, snow, fog, and clear weather, each in day/night) and containing high-quality human-annotated panoptic and semantic ground truth. Each sample includes RGB camera, lidar, radar, and event data, all of which we project onto the RGB plane with the help of the official SDK. We perform all our ablations on this challenging real-world dataset. Due to computational constraints, the hyperparameter ablation in Tab. VI was performed on half of the training data. DeLiVER [29] is a synthetic dataset, featuring 7885 scenes recorded under adverse conditions (cloudy, foggy, night, rainy) and corner-case sensor failures (e.g. motion blur, lidar jitter) with semantic ground truth. Each scene provides RGB camera, lidar, and event camera modalities as well as a pixel-perfect depth map of the scenes. As this depth map is not a direct sensor recording and not realistic to obtain in this quality in any real-world scenario, we argue that one should treat it rather as an additional ground truth. Hence, we use the depth map as the dense supervision for the depth head of our method. As previous methods treat it as an input, we compare to such methods on both scenarios, with (CLDE) and without (CLE) the depth map as an input.
@@ IV-A Comparison with The State of The Art
DGFusion sets the new state of the art (SOTA) in multi-sensor panoptic and semantic segmentation compared to prior methods. As presented in Tab. I, we outperform previous methods on MUSES across all metrics, including a variety of adverse weather and lighting conditions. We set the new SOTA with 61.03% PQ (+1.33%) and improvements are most pronounced in adverse conditions. Compared with the previous SOTA method CAFuser [4], all adverse conditions show a substantial improvement: 1.34% PQ in Fog, 1.63% PQ in Rain and even +2.57% PQ in Snow conditions, compared to a moderate +0.8% PQ in Clear conditions. Further, a Camera + Lidar setup achieves 60.19% PQ, exceeding CAFuser-CL by +1.49%. These findings confirm our hypothesis that fusing both global condition cues and local depth features is crucial for robust performance under adverse conditions.
@@
CMNeXt [29] MiT-B2 50.3 53.0
StitchFusion [20] MiT-B2 50.8 53.4
GeminiFusion [18] MiT-B2 50.5 54.5
CAFuser-CAA [4] Swin-T 51.2 55.2
CAFuser [4] Swin-T 51.3 55.6
@@
As the OneFormer segmentation head can predict both panoptic and semantic segmentation, we also compare DGFusion to SOTA multimodal semantic segmentation methods on the MUSES dataset. Tab. II shows that our method significantly outperforms both camera-only baselines and prior multimodal approaches, achieving 79.5% mIoU (+ 1.0%). Notably, we surpass other fusion-based methods, such as CMNeXt [29] and CAFuser [4], highlighting the effectiveness of our multi-task setup for improving semantic segmentation in adverse-weather driving scenes.
@@
MUSES [2]
OneFormer [17]
CAFuser [4]
@@ IV-C Qualitative Results
In Fig. 3, we present qualitative prediction results on MUSES under various conditions, demonstrating the strengths of our approach compared to existing methods. Under the clear day scenario, we accurately capture the tram in the distance and all four pedestrians on the left sidewalk, whereas other baselines overlook some of these persons. In the second example, we correctly identify wall regions on the side of the bridge in front and detect the bicycle on the left. In the foggy night scene, we are the only method that detects both the cyclist and the bicycle despite the very low visibility. Finally, in the snow daytime scene, DGFusion is the only method that predicts at least one of the two pedestrians on the left sidewalk, though it misses the second person, highlighting a limitation in detecting heavily occluded objects. Our method occasionally confuses similar classes, such as “wall” and “fence” (last row). Nevertheless, these examples highlight how our depth-guided method improves detection under all environmental conditions.
@@ V Conclusion
We have presented DGFusion, a novel depth-guided sensor fusion network for robust semantic perception. Our approach reformulates the multi-sensor fusion task as a multi-task problem by integrating an auxiliary depth head and directly leveraging noisy lidar input as additional supervision. In addition, we propose a depth-guided fusion module that combines local depth tokens with global condition tokens to adaptively guide sensor fusion in any condition. Extensive evaluations on both the real-world MUSES dataset and the synthetic DeLiVER dataset show that DGFusion outperforms previous state-of-the-art methods in panoptic and semantic segmentation, particularly in adverse conditions such as night, fog, rain, and snow. Moreover, our ablation studies confirm the essential contribution of our multi-task loss design and robust depth supervision, demonstrating the value of leveraging noisy lidar depth signals without incurring extra annotation costs. These advancements position DGFusion as a promising solution for enhancing semantic perception in autonomous driving and robotics under challenging environmental scenarios.